INFRARED SPECTRA OF THE 2-CHLOROETHYL RADICAL IN SOLID PARA-HYDROGEN

\underline{JAY C. AMICANGELO}, School of Science, Penn State Erie, Erie, PA 16563; MOHAMMED BA-HOU, BARBARA GOLEC, AND YUAN-PERN LEE, Department of Applied Chemistry and Institute of Molecular Science, National Chiao Tung University, Hsinchu 30010, Taiwan and Institute of Atomic and Molecular Sciences, Academia Sinica, Taipei 10617, Taiwan.

The reaction of chlorine atoms with ethylene and two of its deuterium isotopomers in solid para-hydrogen (p-H_{2}) matrices at 3 K has been studied using infrared spectroscopy. Irradiation at 365 nm of a co-deposited mixture of Cl_{2}, C_{2}H_{4}, and p-H_{2} at 3 K produces a series of new lines in the infrared spectrum. Several of the new lines are readily assigned to the gauche and trans conformers of 1,2-dichloroethane (CH_{2}ClCH_{2}Cl) resulting from the addition of two Cl atoms to C_{2}H_{4}. Of the remaining lines, a strong line at 664 cm^{−1} and three weaker lines at 562, 1070, and 1228 cm^{−1} are concluded to be due to a single carrier based on their behavior upon subsequent annealing to 4.5 K and irradiation at 254 and 214 nm. When the positions and intensities of these lines are compared to the MP2/aug-cc-pVDZ predicted vibrational spectra of the possible species that could result from the addition and abstraction reactions of one Cl atom with C_{2}H_{4}^{a}, the best agreement is found with the 2-chloroethyl radical (·CH_{2}CH_{2}Cl). In order to confirm this assignment, isotopic experiments were performed with C_{2}D_{4} and t-C_{2}H_{2}D_{2} and the corresponding infrared bands due to the deuterium isotopomers of this radical (·CD_{2}CD_{2}Cl and ·CHDCHDCl) have been observed. A final set of experiments were performed following irradiation of the Cl_{2}/C_{2}H_{4}/p-H_{2} mixture at 365 nm, in which the matrix was irradiated with filtered infrared light from a globar source, which has been shown to induce a reaction between isolated Cl atoms and matrix H_{2} to produce HCl and H atoms^{b}. In our experiments, the major products observed were HCl and ethyl chloride (CH_{3}CH_{2}Cl) and the possible mechanism of the formation of ethyl chloride will be discussed.

aP. Brana, B. Menendez, T. Fernandez, and J. A. Sordo, J. Phys. Chem. A \underline{104}, 10842 (2000)

bP. L. Raston and D. T. Anderson, Phys. Chem. Chem. Phys. \underline{8}, 3124 (2006)